Suppose that an account offers a nominal interest rate of 10 %,what is the effective interest rate? if the nominal interest rate is the same, but interest is payable monthly, weekly and daily?


1
Expert's answer
2021-01-21T14:07:44-0500

nominal interest, "r=10\%"

if compounding was monthly, "m=12"

Effective interest rate "=(1+r/m)^{m}-1=(1+0.1/12)^{12}-1=0.1047= 10.47\%"

if the compounding was weekly, "m=52"

Effective interest rate "=(1+r/m)^{m}-1=(1+0.1/52)^{52}-1=0.1051=10.51\%"

if the compounding was daily, "m=365"

Effective interest rate "=(1+r/m)^{m}-1=(1+0.1/365)^{365}-1=0.1052=10.52\%"